The height of an object thrown upward is represented by the equation h=200t-25t^2,where t is the time in seconds and h is the height in meters. What is the maximum height of the ball?
Answer by Alan3354(69443)

The height of an object thrown upward is represented by the equation h=200t-25t^2,where t is the time in seconds and h is the height in meters. What is the maximum height of the ball?
--------------
h(t) = -25t^2 + 200t
The max is the vertex of the parabola.
The vertex is on the Line of Symmetry t = -b/2a = -200/-50 = 4 seconds
h(4) = -25*16 + 200*4
= 400 meters
----------------------